Summary of IM and REC produced today

Everything has been done with FW1 on the wrong position: beamsplitter instead of the dichroic 600-1000

Disturb Basis No modes Type Frames/Movement Push-pull cycle Amplitude file Notes
20120531_154232 KL_v10 200 PP 3 3 600_optimized_3l.fits  
20120531_173955 KL_v11 600 PP 3 1 600_optimized_3l.fits  
20120531_184624 KL_v9 10 PP 3 30 600_optimized_3l.fits  
20120531_185615 KL_v9 100 PP 3 6 600_optimized_3l.fits  
20120531_190646 KL_v9 200 PP 3 3 600_optimized_3l.fits  
20120531_191517 KL_v9 600 PP 3 1 600_optimized_3l.fits  
20120531_213829 KL_v10 600 PP 3 1 600_optimized_3l.fits  
20120531_223651 KL_v11 600 PP 3 1 600_optimized_3l.fits  

IM Basis Disturb Freq Mod Pupils Iteration CL basis CL Rec CL gain Notes
20120531_152048 KL_v10 20120508_000549 600 3 00120530-221500 4 OL OL OL Wrong pupils
20120531_152422 KL_v10 20120508_000549 600 3 00120530-221500 4 KL_v10 20120531_152226 0.05 Wrong pupils
20120531_152919 KL_v10 20120530_222319 600 3 00120530-221500 8 KL_v10 20120531_152629 0.05 Wrong pupils
20120531_154524 KL_v10 20120531_154232 600 3 00120530-221500 8 KL_v10 OL OL Wrong pupils
20120531_155509 KL_v10 20120530_222319 600 3 20120507-230757 48 KL_v10 20120530_234954 0.05 Bad actuator
20120531_170512 KL_v10 20120530_222319 600 3 20120507-230757 48 KL_v10 20120531_160900 0.05 FW1 beamsplitter instead of dichroic
20120531_174112 KL_v11 20120531_173955 990 3 20120507-230757 48 KL_v10 20120531_172208 0.05 Bad 990Hz
20120531_180243 KL_v11 20120531_173955 600 3 20120507-230757 48 KL_v10 20120531_172208 0.05 FW1 beamsplitter instead of dichroic
20120531_185051 KL_v9 20120531_184624 600 3 00120531-163320 4 KL_v10 OL OL FW1 beamsplitter instead of dichroic
20120531_185952 KL_v9 20120531_185615 600 3 00120531-163320 8 KL_v10 20120531_185222 0.05 FW1 beamsplitter instead of dichroic
20120531_190815 KL_v9 20120531_190646 600 3 00120531-163320 8 KL_v10 20120531_190308 0.05 FW1 beamsplitter instead of dichroic
20120531_191642 KL_v9 20120531_191517 600 3 00120531-163320 48 KL_v10 20120531_191123 0.05 FW1 beamsplitter instead of dichroic
20120531_214529 KL_v10 20120531_213829 600 3 00120531-163320 48 KL_v9 20120531_193026 0.05 FW1 beamsplitter instead of dichroic
20120531_223853 KL_v11 20120531_223651 600 3 00120531-163320 48 KL_v9 20120531_193026 0.05 missing file

REC Basis IM Skip Frame Avg Frame Remove TipTilt No modes Cutoff Threshold Notes
20120531_152226 KL_v10 20120531_152048 2 2 Y 10 0 Dont use
20120531_152629 KL_v10 20120531_152422 2 2 Y 10 0 Dont use
20120531_153238 KL_v10 20120531_152919 2 2 Y 100 0 Dont use
20120531_154755 KL_v10 20120531_154524 2 2 Y 100 0 Dont use
20120531_160900 KL_v10 20120531_155509 2 2 Y 400 0 Dont use
20120531_172208 KL_v10 20120531_170512 2 2 Y 400 0 FW1 beamsplitter instead of dichroic
20120531_181634 KL_v11 20120531_180243 2 2 Y 400 0 FW1 beamsplitter instead of dichroic
20120531_185222 KL_v9 20120531_185051 2 2 Y 10 0 FW1 beamsplitter instead of dichroic
20120531_190308 KL_v9 20120531_185952 2 2 Y 100 0 FW1 beamsplitter instead of dichroic
20120531_191123 KL_v9 20120531_190815 2 2 Y 200 0 FW1 beamsplitter instead of dichroic
20120531_193026 KL_v9 20120531_191642 2 2 Y 400 0 FW1 beamsplitter instead of dichroic
20120531_220019 KL_v10 20120531_214529 2 2 Y 400 0 FW1 beamsplitter instead of dichroic

Log

Starting at 06.45 AM or 13.30 UT

The Bayside stage X has some problem with the homing not in the exact zero position.

Check of the position of the AGw probe. Actually it was at the center of the field.

We move the probe: we don't know the password to log on agw1. We use the terminal left open by the TO last night and we send the command: home -u 1 -m 21

Spot is then soon found on IRTC with same global offsets of yesterday

Check of 234954 (KL v10, old pupils, 400 modes)

In the following:
  • old pupils = typically 20120507_230705. Subapertures have been determined with a criterium based on flux measured in day time (small central obstruction)
  • new pupils = pupils derived from 20120507_230705 with a few subaps removed from the center based on actual flux measured during nighttime.

From 20120531_145740 to 20120531_150118 there are some TN acquired with IRTC badly configured. It was meant to check the performance of Rec 20120530_234954 (KL v10, old pupils, 400modes)

Rec 234954 is stable, no particular features are noted in modalplot, some static pattern in the ccd39 frames. It comes from a noisy IM measured during the works on the pier. We decide anyhow to acquire a new IM.

TrackNo Mag Rec bin #mod freq gain mod nph AntiDrift SR filter exp #frames disturb SN notes
20120531_145740 7.4 234954 1 400 990 0.4 0.2 -1.0 3 674 ON -1.0 ? -1 -1 ON NO  
20120531_145757 7.4 234954 1 400 990 0.4 0.2 -1.0 3 669 ON 32.9 EMPTY 10 200 ON NO  
20120531_145849 7.4 234954 1 400 990 0.4 0.2 -1.0 3 675 ON 0.0 J 100 200 ON NO - Unknown IRTC pixelscale
20120531_145936 7.4 234954 1 400 990 0.4 0.2 -1.0 3 678 ON 0.0 J 100 50 ON NO - Unknown IRTC pixelscale
20120531_150054 7.4 234954 1 400 990 0.6 0.4 0.3 3 676 ON 0.0 J 100 50 ON NO - Unknown IRTC pixelscale
20120531_150118 7.4 234954 1 400 990 0.6 0.4 0.3 3 673 ON 0.0 J 100 50 ON NO - Unknown IRTC pixelscale
IDL string: ['20120531_145740','20120531_145757','20120531_145849','20120531_145936','20120531_150054','20120531_150118']

New pupils (00120530-221500)

Note: These pupils are wrong

Note: the GUI to acquire intmat is stucked. Too many disturb files!! It seems related to the NFS filesystem. Typing ls in the directory containing the files to be shown in the drop-down menus seems to be a workaround

We try to compute some REC for KL_v10 with these new pupils but even the 10 modes diverges. It is clear that there is a bug in the way Alfio has derived these pupils from the old ones. We let him fis this problem and we revert to the old pupils

KL_v10 old pupils: second attempt

The 20120530_234954 was created from a noisy IM. We acquire a new IM and Rec (20120531_160900) and we notice during the acquisition that a single actuator on the edge of the mirror has a strange behaviour.

We call Runa at 16:30 to remove the actuator 13 - DSP 39. During this operation we shutdown the ASM and we loose the communication of the derotator angle from AOS to WFS but we did not notice.

TrackNo Mag Rec bin #mod freq gain mod nph AntiDrift SR filter exp #frames disturb SN notes
20120531_161828 8.0 160900 1 400 990 0.4 0.5 0.4 3 376 ON -1.0 ? -1 -1 ON NO  
20120531_161901 8.0 160900 1 400 990 0.4 0.5 0.4 3 373 ON 32.6 J 100 50 ON NO  
20120531_161923 8.0 160900 1 400 990 0.4 0.5 0.4 3 374 ON 31.8 J 100 50 ON NO  
20120531_161952 8.0 160900 1 400 990 0.4 0.5 0.4 3 373 ON 31.6 J 100 50 ON NO  
20120531_162018 8.0 160900 1 400 990 0.4 0.5 0.4 3 377 ON 32.1 J 100 50 ON NO  
20120531_163735 8.2 160900 1 400 990 0.6 0.4 0.4 3 310 ON 37.0 J 100 50 ON NO  
20120531_163915 8.2 160900 1 400 990 0.6 0.4 0.4 3 312 ON 37.8 J 100 50 ON NO  
20120531_163934 8.2 160900 1 400 990 0.6 0.4 0.4 3 311 ON 34.8 J 100 50 ON NO  

modalplotSr30.jpg

Screenshot_120531_0.png

The pattern on the border is due to a wrong position of the rotator 2deg. Once corrected the effect disappeared. The telescope rotator was slowly moving. We stopped it.

KL_v10 old pupils: third and final attempt

We measure a new IM KL_V10 old pupils (20120531_170512) and REC 400 modes (20120531_172208)

TrackNo Mag Rec bin #mod freq gain mod nph AntiDrift SR filter exp #frames disturb SN notes
20120531_173104 7.6 172208 1 400 990 0.4 0.4 0.4 3 531 ON -1.0 ? -1 -1 ON NO  
20120531_173120 7.6 172208 1 400 990 0.4 0.4 0.4 3 532 ON 33.9 J 100 50 ON NO IRTC dark +1h old!
20120531_173248 7.6 172208 1 400 990 0.4 0.4 0.4 3 532 ON 27.5 J 100 50 ON NO IRTC dark +1h old!
20120531_173302 7.6 172208 1 400 990 0.4 0.4 0.4 3 532 ON 24.1 J 100 50 ON NO IRTC dark +1h old!
20120531_173317 7.6 172208 1 400 990 0.4 0.4 0.4 3 530 ON 29.5 J 100 50 ON NO IRTC dark +1h old!
20120531_173330 7.6 172208 1 400 990 0.4 0.4 0.4 3 528 ON 29.9 J 100 50 ON NO IRTC dark +1h old!

modalplotKL_V10_400modes_new.jpg

KL_v11 old pupils

ScreenshotKL_V11.png

TrackNo RefStar Mag El Wind DIMM/OL Rec bin #mod freq gain mod nph AntiDrift SR filter exp #frames disturb SN notes
20120531_182407 8.1 181634 1 400 990 0.6 0.3 -1.0 3 355 ON -1.0 ? -1 -1 ON NO  
20120531_182431 8.1 181634 1 400 990 0.6 0.3 -1.0 3 353 ON 25.0 J 100 50 ON NO  
20120531_182450 8.1 181634 1 400 990 0.6 0.3 -1.0 3 354 ON 37.8 J 100 50 ON NO  
20120531_182504 8.1 181634 1 400 990 0.6 0.3 -1.0 3 355 ON 38.2 J 100 50 ON NO  
20120531_182517 8.1 181634 1 400 990 0.6 0.3 -1.0 3 356 ON 37.7 J 100 50 ON NO  
20120531_182804 8.2 181634 1 400 990 0.4 0.5 0.4 3 310 ON 29.1 J 100 50 ON NO  
20120531_182816 8.2 181634 1 400 990 0.4 0.5 0.4 3 310 ON 29.1 J 100 50 ON NO  
20120531_182829 8.2 181634 1 400 990 0.4 0.5 0.4 3 310 ON 28.3 J 100 50 ON NO  
20120531_182842 8.2 181634 1 400 990 0.4 0.5 0.4 3 311 ON 19.4 J 100 50 ON NO  

ModalPlotKL_v11_single.jpg ModalPlotKL_v11_multi.jpg

KL_v9 new pupils (00120531_163320)

Alfio fixed the bug and produced new pupils 00120531_163320

We save 10, 100, 200 and 600 INTMAT iteratively: 10 in open loop, 100 using the RECMAT 10 modes etc...

TrackNo Mag Rec bin #mod freq gain mod nph AntiDrift SR filter exp #frames disturb SN notes
20120531_193612 7.6 193026 1 400 990 0.5 0.5 0.3 3 573 ON -1.0 ? -1 -1 ON NO  
20120531_193627 7.5 193026 1 400 990 0.5 0.5 0.3 3 579 ON 33.5 J 100 50 ON NO  
20120531_193641 7.5 193026 1 400 990 0.5 0.5 0.3 3 577 ON 31.5 J 100 50 ON NO  
20120531_193654 7.5 193026 1 400 990 0.5 0.5 0.3 3 577 ON 27.6 J 100 50 ON NO  
20120531_193707 7.5 193026 1 400 990 0.5 0.5 0.3 3 580 ON 28.0 J 100 50 ON NO  
20120531_193857 7.5 193026 1 400 990 0.5 0.5 0.3 3 580 ON 28.1 J 100 50 ON NO  
20120531_193913 7.5 193026 1 400 990 0.5 0.5 0.3 3 580 ON 33.0 J 100 50 ON NO  
20120531_193926 7.5 193026 1 400 990 0.5 0.5 0.3 3 584 ON 33.7 J 100 50 ON NO  
20120531_193940 7.5 193026 1 400 990 0.5 0.5 0.3 3 584 ON 30.1 J 100 50 ON NO  
20120531_201520 7.5 193026 1 400 990 0.5 0.5 0.3 3 589 ON 24.9 J 50 200 ON NO IRTC dark +1h old!
20120531_201552 7.5 193026 1 400 990 0.5 0.5 0.3 3 585 ON 10.0 J 50 20 ON NO IRTC dark +1h old!
20120531_201607 7.5 193026 1 400 990 0.5 0.5 0.3 3 588 ON 32.1 J 50 20 ON NO  
20120531_201636 7.5 193026 1 400 990 0.5 0.5 0.3 3 590 ON 23.0 J 50 20 ON NO  
20120531_201912 7.5 193026 1 400 990 0.5 0.5 0.3 3 591 ON 20.5 J 50 20 ON NO  
20120531_201927 7.5 193026 1 400 990 0.5 0.5 0.3 3 591 ON 28.6 J 50 20 ON NO  
20120531_201942 7.5 193026 1 400 990 0.5 0.5 0.3 3 592 ON 26.3 J 50 20 ON NO  
20120531_202023 7.5 193026 1 400 990 0.5 0.5 0.3 3 588 ON 17.3 J 50 20 ON NO  
20120531_203721 7.5 193026 1 400 990 0.5 0.5 0.3 3 590 ON 29.4 J 50 20 ON NO  
20120531_204109 7.5 193026 1 400 990 0.5 0.5 0.3 3 591 ON 24.2 J 50 20 ON NO  
20120531_204338 7.5 193026 1 400 990 0.5 0.5 0.3 3 593 ON 31.2 J 50 20 ON NO  
20120531_204859 7.5 193026 1 400 990 0.6 0.4 0.4 3 616 ON 32.9 J 50 20 ON NO  

ModalPlotKL_v9NewPupils.jpg

Call with Simone and Alfio. These SR are lame: in 20120203_182114 et simili they had SR=45% in J band.

The following measurment were taken during the call. The difference between the sides is sometimes suspiciously high 0.6px.

TrackNo RefStar Mag Rec bin #mod freq gain mod nph AntiDrift SR filter exp #frames disturb SN notes
20120531_213414 DuringCall 7.4 193026 1 400 990 0.5 0.3 0.2 3 674 ON 35.0 J 50 200 OFF NO  
20120531_213442 DuringCall 7.4 193026 1 400 990 0.5 0.3 0.2 3 674 ON 37.7 J 50 200 OFF NO  
20120531_213507 DuringCall 7.4 193026 1 400 990 0.5 0.3 0.2 3 673 ON 30.7 J 50 200 OFF NO  
20120531_213542 DuringCall 7.4 193026 1 400 990 0.5 0.3 0.2 3 676 ON 36.3 J 50 200 OFF NO  

It seems we are also out of focus on the IRTC (39.9mm). They had z=39.6mm with cube, 10mas/px, J band.

ModalPlotKL_v9_disturbOff.jpg

KL_v10 new pupils

We measure the reconstructor 220019

Run autofocus. Very strange: best autofocus position is 40.2 instead of 36.60. Repeated twice

It seems that the measurements confim this! But we are still far away from SR=45%

TrackNo RefStar Mag Rec bin #mod freq gain mod nph AntiDrift SR filter exp #frames disturb SN notes
20120531_222239 RetroAfterCall 7.9 220019 1 400 990 0.6 0.5 0.4 3 406 ON 36.3 J 50 200 ON NO z=40.17
20120531_222330 RetroAfterCall 7.9 220019 1 400 990 0.6 0.5 0.4 3 411 ON 35.5 J 50 200 ON NO z=40.17
20120531_222355 RetroAfterCall 7.9 220019 1 400 990 0.6 0.5 0.4 3 410 ON 34.3 J 50 200 ON NO z=40.17
20120531_222425 RetroAfterCall 7.9 220019 1 400 990 0.6 0.5 0.4 3 409 ON 36.0 J 50 200 ON NO z=40.17
20120531_222740 RetroAfterCall 7.8 220019 1 400 990 0.6 0.4 0.5 3 472 ON 32.3 J 50 200 ON NO z=40.17 + Autogain
20120531_222805 RetroAfterCall 7.8 220019 1 400 990 0.6 0.4 0.5 3 472 ON 32.5 J 50 200 ON NO z=40.17 + Autogain
20120531_222831 RetroAfterCall 7.8 220019 1 400 990 0.6 0.4 0.5 3 473 ON 33.2 J 50 200 ON NO z=40.17 + Autogain
20120531_222918 RetroAfterCall 7.8 220019 1 400 990 0.6 0.4 0.5 3 460 ON 26.8 J 50 200 ON NO z=39.6
20120531_223036 RetroAfterCall 7.8 220019 1 400 990 0.6 0.4 0.5 3 472 ON 27.4 J 50 200 ON NO z=39.6
20120531_223110 RetroAfterCall 7.8 220019 1 400 990 0.6 0.4 0.5 3 473 ON 26.3 J 50 200 ON NO z=39.6

ModalPlotKL_v10_newpupils.jpg

KL_v11 new pupils

We measure the IM 223853. Error in analyzing the IM Alfio???

Why are SR so low?

The strange thing is that the modalplots obtained today seems totally compatible with the ones of 20120203, still the PSF is visibly worse and SR too of course.

Test without disturbance

TrackNo Mag Rec bin #mod freq gain mod nph AntiDrift SR filter exp #frames disturb SN notes
20120531_230411 7.8 193026 1 400 990 0.5 0.2 -1.0 3 462 ON -1.0 ? -1 -1 OFF NO z=36.6
20120531_230428 7.8 193026 1 400 990 0.5 0.2 -1.0 3 462 ON 37.9 J 50 100 OFF NO z=36.6
20120531_230454 7.8 193026 1 400 990 0.5 0.2 -1.0 3 461 ON 34.3 J 50 100 OFF NO z=36.6
20120531_230508 7.8 193026 1 400 990 0.5 0.2 -1.0 3 461 ON 28.0 J 50 100 OFF NO z=36.6
20120531_230527 7.8 193026 1 400 990 0.5 0.2 -1.0 3 462 ON 34.0 J 50 100 OFF NO z=36.6
20120531_230653 7.8 193026 1 400 990 0.5 0.2 -1.0 3 463 ON -1.0 ? -1 -1 OFF NO z=36.6
20120531_230708 7.8 193026 1 400 990 0.5 0.2 -1.0 3 464 ON 29.0 J 50 100 OFF NO z=36.6
20120531_230725 7.8 193026 1 400 990 0.5 0.2 -1.0 3 463 ON 32.0 J 50 100 OFF NO z=36.6
20120531_230740 7.8 193026 1 400 990 0.5 0.2 -1.0 3 463 ON 37.6 J 50 100 OFF NO z=36.6
20120531_230757 7.8 193026 1 400 990 0.5 0.2 -1.0 3 462 ON 28.5 J 50 100 OFF NO z=36.6
20120531_230838 7.8 193026 1 400 990 0.5 0.2 -1.0 3 466 ON 41.0 J 50 100 OFF NO z=40.2
20120531_230902 7.8 193026 1 400 990 0.5 0.2 -1.0 3 466 ON 39.7 J 50 100 OFF NO z=40.2
20120531_230924 7.8 193026 1 400 990 0.5 0.2 -1.0 3 467 ON 37.1 J 50 100 OFF NO z=40.2
20120531_231015 7.8 193026 1 400 990 0.5 0.2 -1.0 3 465 ON 257.0 J 50 100 OFF NO z=40.2 IRTC T > 260K

Now we decide to test the old 500modes rec of February. Change pupil with change_binning

Without disturb

TrackNo RefStar Mag Rec bin #mod freq gain mod nph AntiDrift SR filter exp #frames disturb SN notes
20120531_233405 KL_v9 Feb 7.4 181326 1 500 990 0.4 0.2 -1.0 3 649 ON 35.8 J 50 50 OFF NO  
20120531_233423 KL_v9 Feb 7.4 181326 1 500 990 0.4 0.2 -1.0 3 649 ON 21.3 J 50 100 OFF NO  
20120531_233438 KL_v9 Feb 7.4 181326 1 500 990 0.4 0.2 -1.0 3 649 ON 31.0 J 50 100 OFF NO  

And with disturbance

TrackNo RefStar Mag Rec bin #mod freq gain mod nph AntiDrift SR filter exp #frames disturb SN notes
20120531_233759 KL_v9 Feb 7.4 181326 1 500 990 0.5 0.5 0.4 3 654 ON 22.2 J 50 100 ON NO  
20120531_233817 KL_v9 Feb 7.4 181326 1 500 990 0.5 0.5 0.4 3 655 ON 31.5 J 50 100 ON NO  
20120531_233833 KL_v9 Feb 7.4 181326 1 500 990 0.5 0.5 0.4 3 649 ON 24.6 J 50 100 ON NO  

SR are not ok. We are not reproducing the SR=45 of Feb

Carmelo supposes IRTC is shaking. Rotation of 45 degrees of the Rotator: from 286 to 331.

TrackNo RefStar Mag Rec bin #mod freq gain mod nph AntiDrift SR filter exp #frames disturb SN notes
20120531_235617 KL_v9 Feb 7.4 181326 1 500 990 0.7 0.3 0.3 3 652 ON 32.5 J 50 100 ON NO  
20120531_235636 KL_v9 Feb 7.4 181326 1 500 990 0.7 0.3 0.3 3 650 ON 45.1 J 50 100 ON NO  
20120531_235653 KL_v9 Feb 7.4 181326 1 500 990 0.7 0.3 0.3 3 649 ON 42.8 J 50 100 ON NO  
20120601_000103 KL_v9 Feb 7.4 181326 1 500 990 0.6 0.4 0.3 3 649 ON 43.5 J 50 100 ON NO  
20120601_000123 KL_v9 Feb 7.4 181326 1 500 990 0.6 0.4 0.3 3 644 ON 43.3 J 50 100 ON NO  
20120601_000147 KL_v9 Feb 7.4 181326 1 500 990 0.6 0.4 0.3 3 648 ON 45.5 J 50 100 ON NO  
20120601_000203 KL_v9 Feb 7.4 181326 1 500 990 0.6 0.4 0.3 3 653 ON 44.1 J 50 100 ON NO  

Wow: it worked (?)

Let's try with no disturb

TrackNo RefStar Mag Rec bin #mod freq gain mod nph AntiDrift SR filter exp #frames disturb SN notes
20120601_000834 KL_v9 Feb 7.4 181326 1 500 990 0.5 0.2 0.2 3 643 ON 48.6 J 50 100 OFF NO  
20120601_000856 KL_v9 Feb 7.4 181326 1 500 990 0.5 0.2 0.2 3 640 ON 46.8 J 50 100 OFF NO  
20120601_000911 KL_v9 Feb 7.4 181326 1 500 990 0.5 0.2 0.2 3 642 ON 49.1 J 50 100 OFF NO  
20120601_000927 KL_v9 Feb 7.4 181326 1 500 990 0.5 0.2 0.2 3 641 ON 50.9 J 50 100 OFF NO  

Stil high, even if it is not 70%

So maybe now it is the good moment. Let's go back to the new reconstructors measured today

TrackNo RefStar Mag Rec bin #mod freq gain mod nph AntiDrift SR filter exp #frames disturb SN notes
20120601_001409 KL_v9 New Pupil 7.4 193026 1 400 990 0.6 0.4 0.3 3 663 ON 43.4 J 50 100 ON NO - SR may be in error
20120601_001433 KL_v9 New Pupil 7.4 193026 1 400 990 0.6 0.4 0.3 3 655 ON 44.0 J 50 100 ON NO  
20120601_001447 KL_v9 New Pupil 7.4 193026 1 400 990 0.6 0.4 0.3 3 660 ON 40.9 J 50 100 ON NO - SR may be in error
20120601_001501 KL_v9 New Pupil 7.4 193026 1 400 990 0.6 0.4 0.3 3 659 ON 38.1 J 50 100 ON NO - SR may be in error
20120601_001516 KL_v9 New Pupil 7.4 193026 1 400 990 0.6 0.4 0.3 3 660 ON 44.8 J 50 100 ON NO - SR may be in error
20120601_001627 KL_v9 New Pupil 7.4 193026 1 400 990 0.6 0.4 0.3 3 661 ON 42.3 J 50 100 ON NO  
20120601_001642 KL_v9 New Pupil 7.4 193026 1 400 990 0.6 0.4 0.3 3 660 ON 44.5 J 50 100 ON NO - SR may be in error
20120601_001657 KL_v9 New Pupil 7.4 193026 1 400 990 0.6 0.4 0.3 3 662 ON 43.0 J 50 100 ON NO - SR may be in error

Very good: we are finally able to reproduce the famous 45% in J

Let's try to move a little bit in the field

TrackNo RefStar Mag Rec bin #mod freq gain mod nph AntiDrift SR filter exp #frames disturb SN notes
20120601_001948 KL_v9 New Pupil 7.4 193026 1 400 990 0.6 0.4 0.3 3 669 ON 36.5 J 50 100 ON NO  
20120601_002009 KL_v9 New Pupil 7.4 193026 1 400 990 0.6 0.4 0.3 3 669 ON 35.7 J 50 100 ON NO  
20120601_002025 KL_v9 New Pupil 7.4 193026 1 400 990 0.6 0.4 0.3 3 669 ON 35.2 J 50 100 ON NO  
20120601_002040 KL_v9 New Pupil 7.4 193026 1 400 990 0.6 0.4 0.3 3 669 ON 34.1 J 50 100 ON NO  

Oh... That's bad!! What happened? We didn't understand anything...

Again an offset

TrackNo RefStar Mag Rec bin #mod freq gain mod nph AntiDrift SR filter exp #frames disturb SN notes
20120601_002227 KL_v9 New Pupil 7.4 193026 1 400 990 0.6 0.4 0.3 3 669 ON 35.2 J 50 100 ON NO  
20120601_002248 KL_v9 New Pupil 7.4 193026 1 400 990 0.6 0.4 0.3 3 667 ON 35.4 J 50 100 ON NO  
20120601_002303 KL_v9 New Pupil 7.4 193026 1 400 990 0.6 0.4 0.3 3 666 ON 35.2 J 50 100 ON NO  
20120601_002318 KL_v9 New Pupil 7.4 193026 1 400 990 0.6 0.4 0.3 3 667 ON 34.4 J 50 100 ON NO  
20120601_002456 KL_v9 New Pupil 7.4 193026 1 400 990 0.5 0.4 0.3 3 669 ON 34.4 J 50 100 ON NO  
20120601_002514 KL_v9 New Pupil 7.4 193026 1 400 990 0.5 0.4 0.3 3 667 ON 32.6 J 50 100 ON NO  
20120601_002528 KL_v9 New Pupil 7.4 193026 1 400 990 0.5 0.4 0.3 3 666 ON 33.2 J 50 100 ON NO  
20120601_002543 KL_v9 New Pupil 7.4 193026 1 400 990 0.5 0.4 0.3 3 668 ON 32.8 J 50 100 ON NO  

Topic attachments
I Attachment Action Size Date Who Comment
ModalPlotKL_v10_newpupils.jpgjpg ModalPlotKL_v10_newpupils.jpg manage 78 K 31 May 2012 - 22:41 CarmeloArcidiacono KL_v10 new pupils
ModalPlotKL_v11_multi.jpgjpg ModalPlotKL_v11_multi.jpg manage 83 K 31 May 2012 - 18:40 CarmeloArcidiacono all meas KL_v11
ModalPlotKL_v11_single.jpgjpg ModalPlotKL_v11_single.jpg manage 62 K 31 May 2012 - 18:40 CarmeloArcidiacono one meas KL_v11
ModalPlotKL_v9NewPupils.jpgjpg ModalPlotKL_v9NewPupils.jpg manage 74 K 31 May 2012 - 21:30 CarmeloArcidiacono KL_v9 New Pupils
ModalPlotKL_v9_disturbOff.jpgjpg ModalPlotKL_v9_disturbOff.jpg manage 67 K 31 May 2012 - 21:43 CarmeloArcidiacono KL_v9 New Pupils Disturb off
Screenshot400.pngpng Screenshot400.png manage 892 K 31 May 2012 - 17:28 CarmeloArcidiacono  
ScreenshotKL_V11.pngpng ScreenshotKL_V11.png manage 856 K 31 May 2012 - 18:20 CarmeloArcidiacono KL_V11
Screenshot_120531_0.pngpng Screenshot_120531_0.png manage 815 K 31 May 2012 - 16:55 CarmeloArcidiacono KL_V10 400 modes
modalplotKL_V10_400modes_new.jpgjpg modalplotKL_V10_400modes_new.jpg manage 66 K 31 May 2012 - 18:22 CarmeloArcidiacono modalplot KL_v10
modalplotSr30.jpgjpg modalplotSr30.jpg manage 61 K 31 May 2012 - 16:55 CarmeloArcidiacono KL_V10 400 modes
Topic revision: r10 - 11 Sep 2012, EnricoPinna
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback